OpenStudy (drakedoesmc):

Ms. Wong is redecorating her office. She has a choice of 7 colors of paint, 4 kinds of curtains, 3 colors of carpet, and 2 styles of furniture. How many different ways are there to redecorate if she chooses two different colors of paint, one kind of curtain, one color of carpet, and one style of furniture? A. 168 ways B. 1,008 ways C. 1,176 ways D. 23 ways

OpenStudy (nurali):

4 types of curtain, 3 colors carpet and 2 of furniture number of ways = 21 * 4 * 3 * 2 = 21 * 24 = 504* 2=1008 Ways